# Fine-Grained Alignment and Noise Refinement for Compositional Text-to-Image Generation

### v1

@@ 5 Conclusion
Our work mitigates compositional challenges through targeted objectives. These objectives are designed to complement each other and work collaboratively to structure the image scene based. Our fine-grained initial noise refinement framework employs a feedback-driven system with a verifier that provides fine-grained feedback, enabling precise adjustments to the initial noise by targeting only the faulty attention maps while preserving the correct ones.
Supplementary Material
6 PRELIMINARIES
Stable Diffusion. We apply our evaluation and refinement method on the Stable Diffusion model [34]. In SD, an encoder $\mathcal{E}$ is trained to map an image $x\in\mathcal{X}$ to a latent code $z=\mathcal{E}(x)$. The latent code $z$ is then given to a decoder model to reconstruct the input image such that $\mathcal{D}(\mathcal{E}(x))\approx x$. After training the autoencoder, a denoising diffusion probabilistic model (DDPM) [16] is trained over the latent space of the autoencoder. The UNet [35] denoising model $\epsilon_{\theta}$ learns to denoise an input latent $z_{t}$ at each timestep $t$ (where $z_{t}$ results from adding $\epsilon$ noise gradually over $t$ timesteps to the original latent $z_{0}$ during training). The denoising objective which intends to learn $\epsilon_{\theta}$ in order to predict the noise $\epsilon$ is given by:
The denoising process of Stable Diffusion is conditioned on the embedding of text information $L(\mathcal{P})$. In Stable Diffusion, this embedding is the output of the CLIP [31] text encoder. At inference time, a random latent $z_{t}$ is sampled from $\mathcal{N}(0,I)$, and the trained UNet denoising model $\epsilon_{\theta}$ outputs a denoised latent $z_{0}$. We obtain the reconstructed image by inputting the $z_{0}$ to the decoder $\mathcal{D}$.
The cross-attention map has the shape of $h\times w\times l$ where $h$ and $w$ are feature map resolution and $l$ is the number of text tokens. We set feature map resolution to $16\times 16$ as it empirically contains the most semantically meaningful attention maps [13].
7 Related works
Text-to-Image Generative Models. Before the emergence of diffusion-based models, various research directions aimed to achieve realistic and high-quality image generation, both conditional and unconditional, through approaches like generative adversarial networks (GANs) [20, 41, 44, 48, 53, 47] and autoregressive models [5, 8, 32, 46]. Recently, diffusion-based models have shown remarkable performance in conditional text-to-image generation [34, 29, 36, 33, 4]. While having a good performance, they often have alignment problems when generating images. We discuss previous studies on alignment problems under three categories.
Training-Based Compositional Improvement. A group of studies have focused on overcoming compositional challenges in text-conditional diffusion-based models to improve alignment between text prompts and generated images. Some approaches address this through training-based methods [22, 43, 27, 49, 19, 51, 9]. For instance, T2I-Adapter [27] and ControlNet [49] target the control of semantic structures by specifying high-level features. ReCo [43] refines spatial awareness using adapters on top of the diffusion models, while GLIGEN [22] integrates grounding information into newly trainable layers. Composer [19] decomposes images into key factors, training a diffusion model with these factors as conditions to recompose the input. Alternatively, Itercomp [51] utilizes iterative feedback learning to enhance the compositional generation, and ReNo [9] adopts reward-based noise optimization for improved alignment. While being effective, adopting these methods comes with the cost of additional training time and resources.
Training-Free Compositional Improvement. To mitigate compositional problems, some methods leverage training-free methods using latent or attention maps [25, 10, 6, 2, 23, 26, 12, 37]. For instance, Composable Diffusion [25] computes separate denoising latent for each statement and combines them with a score function, or Structure Diffusion [10] manipulates the cross-attention maps guided by hierarchically extracted structures from text. Attend-and-Excite introduced a novel loss based on cross-attention maps is utilized to optimize noise latent, aiming to reduce missing objects. Similarly, A-STAR [2] optimizes noise latent by employing segregation and retention losses to minimize cross-attention overlap between different concepts, leading the objects not to be overlooked while preserving cross-attention information across all concepts. Also, Divide-and-Bind [23] addresses the problem of entity missing by leveraging an attendance objective. Additionally, it enhances attribute association through the incorporation of a binding objective. Similarly, CONFORM [26] addresses entity missing and attribute binding issues by employing a contrastive objective. This approach brings each entity and its attribute’s attention map closer together while disentangling the attention maps of different entities and their respective attributes. In another approach [12], the initial noise is optimized to ensure it lies within a valid space for generating images aligned with the input prompt. Another approach involves utilizing evaluation feedback to refine the image after its generation. The Evaluate-and-Refine method [37] employs iterative VQA feedback to adjust the weighting of CLIP’s phrase embeddings. Additionally, it integrates the Attend-and-Excite method to address the issue of entity missing effectively. However, these approaches are restricted in handling compositional problems such as spatial relationships, entity missing, and attribute binding in an integrated manner because they do not leverage detailed textual information properly. Furthermore, utilizing detailed textual information enables us to arrange the scene more effectively.
Leveraging LLMs for Accurate Alignment in Text-to-Image Models. Recent works have leveraged layout-based approaches to enhance compositional alignment in images generated by text-to-image models, using spatial layouts or bounding boxes [24, 50, 22, 39, 52]. Although these methods enhance spatial awareness, they face challenges in achieving realistic image generation, particularly in creating non-spatial relationships and maintaining aesthetic quality [50, 51]. Other studies utilizing LLMs aim to break down the generation of an aligned image by converting the prompt into simpler subtasks, with the planning of these subtasks throughout the generation process [24, 45, 21, 42]. However, these models often struggle to produce accurate results due to the inherent complexity of LLM outputs [42, 51]. Additionally, since these methods also extract spatial layouts or bounding boxes, they encounter challenges similar to those faced by layout-based methods.
8 Verifier Adaptation
@@
We modify the DA-Score [37] to provide fine-grained feedback regarding the types of problems present in the image. To do so, we categorize questions by the type of issue associated with each entity. At first, similar to the DA-Score, we extract a diverse set of questions covering all entities. Then, we remove extra details from questions with one entity to create coarse-grained questions to address entity missing. Simultaneously, we categorize each question into one of the attribute binding, or spatial relation problems using LLMs [1]. After categorization, we assign a score to each question using a VQA model like DA-Score. For each entity, we aggregate question scores in every category to obtain the final fine-grained feedback. We then use the feedback provided by our evaluator for these questions to assign three scores to each entity, assessing their generation quality in each category.
9 Experiments
9.1 Benchmarks
@@
HRS-Bench Benchmark. This benchmark [3] evaluates text-to-image models across 13 skills, grouped into five key areas: accuracy, robustness, generalization, fairness, and bias. Spanning 50 scenarios, including fashion, animals, transportation, food, and clothing, it offers a comprehensive evaluation framework. To demonstrate our model’s ability to generate multi-entity images, we utilized the 3-entity prompts from its color category.
9.2 Metrics
9.2.1 Alignment Metrics
We employed various metrics to evaluate our method and compare it with other methods. The most commonly used metric for assessing Text-to-Image models is the CLIP-Score [14], which measures the similarity between the image and text embeddings generated by the CLIP model. The TIFA-Score [17] utilizes a language model to generate multiple question-answer pairs from the input prompt, which are then filtered using a QA model. A VQA model answers these visual questions based on the generated image, and the correctness of the answers is evaluated. Additionally, the DA-Score [37] breaks the prompt into a set of disjoint assertions and evaluates their alignment with the generated image directly using a VQA model. VQA models often struggle to accurately interpret spatial relationships. We employ position-based metrics as a more dependable approach for evaluating generative models with respect to spatial relationships. We utilize the VISOR Score [11], designed to assess the spatial relationships of entities. This score is derived by extracting entities using an object detector and calculating the positions of object bounding box centroids relative to one another. The other automated metric we used was the reward model introduced in [40]. This reward model is trained on human preference comparisons to predict the alignment of generated images with textual prompts, providing an evaluation metric for the quality of text-to-image generation based on human preferences.
9.2.2 Human Evaluation
@@
9.3 More ablation Study
@@
9.3.1 Fine-grained Refinement Method Ablation Study
@@
9.3.2 Diversity & Reality
To evaluate the realism and diversity of generated images and compare models, we used the Density and Coverage metrics proposed by [28]. Density measures fidelity by assessing how many real data points lie within the neighborhood of each generated sample, reflecting how well the generated data aligns with the real distribution. Coverage quantifies diversity by determining the proportion of real data points that have at least one generated sample in their neighborhood, indicating how effectively the model captures the variability of real data.
9.4 Additional Quantitative Results
9.4.1 Attribute Binding Loss Comparison
We evaluated three different losses for attribute binding. As shown in Table 8, this analysis demonstrates the effectiveness of our IoU-based loss for addressing the attribute binding problem, outperforming other losses employed in previous studies [23].
@@ 9.4.2 Additional Metrics
Similarly, as demonstrated in Table 7, our method outperforms all other models on the CLIP-Score metric for the HRS benchmark and most categories of the T2I-CompBench benchmark, while achieving comparable performance in the ”Complex” and ”Texture” categories. On the other hand, this metric struggles to underscore the differences between various methods, as the score variations among different models are smaller compared to other metrics. This limitation arises from the inherent weaknesses of the CLIP [31] model in compositional understanding, primarily due to its training methodology and the scoring mechanism, which overlooks several aspects of compositionally.
@@
Table 9: Evaluation of image realism and diversity.
9.5 Additional Qualitative Results
Figure 10 presents a side-by-side comparison of our method with state-of-the-art models, focusing exclusively on spatial relations, while Figure 13 compares our approach with other baselines across all categories. These figures highlight common failure cases in existing models that our method effectively addresses, such as entity missing in ”a blue giraffe and a brown vase” and attribute binding errors like ”a yellow banana and a red monkey.” Additionally, our method successfully handles complex spatial relationships where other models struggle, as demonstrated in ”a brown teddy bear holding a blue cup sits on the sofa.”
Comparison of our models with fine-tuned models We compare our models with GORS [18], a method that fine-tunes Stable Diffusion v2 using generated images closely aligned with compositional prompts. In this approach, the fine-tuning loss is weighted by a reward based on alignment scores. As demonstrated in Fig. 12, our models outperform GORS on the T2I-CompBench [18].
@@ 9.5 Additional Qualitative Results
Figure 11: Visualization of the final cross-attention maps. Compare the final cross-attention maps of Stable Diffusion v1.4 and our fine-grained initial noise refinement.
@@
9.6 Attention maps as object detection
We extracted object boxes from attention maps to qualitatively assess the generated images. The details for bounding box extraction are provided in section 10. Fig. 8 illustrated some more visualizations of entity boxes from our generated images. Extracted boxes show great alignment with the location where the image is actually generated, showcasing the proper appearance of multiple entities.
@@ 9.6 Attention maps as object detection
10 Implementation Details
@@
Given a caption: Specify the spatial location among objects in the caption as (object, spatial relation, object). If there is not any spatial relation in the caption, just say ”No spatial” Example: Caption: ”the striped rug was on top of the tiled floor” 1. (rug, on top of, floor) Example: Caption: ”a couple is enjoying a picnic in the park” 1. (couple, in, park) Example: Caption: ”a blue scooter is parked near a curb in front of a green vintage car” 1. (scooter, near, curb) 2. (scooter, in front of, car) Example: Caption: ”the airplane is flying above the clouds.” 1. (airplane, above, clouds) Example: Caption: ”a bird on the left of a clock” 1. (bird, on the left of, clock) Example: Caption: ”the black phone was resting on the silver charger” 1. (phone, on, charger) Example: Caption: ”the bear is near a lake” 1. (bear, near, lake) Example: Caption: ”a mouse on side of a bag” 1. (mouse, on side of, bag) Example: Caption: ”a small white kitchen with brown wood floor” No spatial Answer as concisely as possible.
@@
Given a caption: Specify the adjective and its object in the caption as (adjective, object) If there is not any adjective in the caption, just say ”No adjective” Example: Caption: ”a brown dog fetching a frisbee in a green park.” 1. (brown, dog) 2. (green, park) Example: Caption: ”a long necklace and a short earring” 1. (long, necklace) 2. (short, earring) Example: Caption: ”a soft pillow is on top of the rocking chair” 1. (soft, pillow) 2. (rocking, chair) Example: Caption: ”a metallic bracelet and a leather hat” 1. (metallic, bracelet) 2. (leather, hat) Example: Caption: ”a person is wearing a hat and sunglasses while fishing” No adjective Example: Caption: ”a desk on the right of a horse” No adjective Example: Caption: ”the shiny silver watch lay next to the smooth black leather wallet” 1. (shiny, watch) 2. (silver, watch) 3. (smooth, wallet) 4. (black, wallet) 5. (leather, wallet) Answer as concisely as possible.

### v2

@@ 7 Conclusion
Our work improves text-to-image generative models by addressing compositional challenges through targeted objectives—including entity missing, entity mixing, attribute binding, and spatial relation losses—that enable precise, fine-grained noise adjustments. These interpretable objectives are designed to complement each other and work collaboratively to structure the image scene based on the entities and their relationships mentioned in the prompt. Additionally, our fine-grained initial noise refinement framework employs a feedback-driven system with a verifier that provides fine-grained feedback to separate entities into proper and faulty ones, enabling precise adjustments to the initial noise by targeting only the faulty attention maps while preserving the correct ones. This process iteratively adjusts the initial noise and effectively addresses generation issues with multiple entities.
8 Limitations
Our proposed training-free method offers a novel and efficient solution for addressing compositional generation failure modes in T2I models. Nevertheless, it has several limitations that we elaborate on in this section.
Struggle with long and complex prompts. As shown in (Zhang et al., 2024a; Liu et al., 2024), diffusion models that rely on CLIP-based text encoders often struggle with long captions. This is because the CLIP text encoder has limited ability to understand and represent the compositional structure of complex or lengthy prompts. Since our method does not modify the underlying text encoder, it inherits this limitation and may similarly face challenges when processing long or complex prompts.
Non-Optimal Component Design. Our method introduces an innovative approach by designing a cost function that integrates objectives addressing multiple types of constraints simultaneously. However, our primary goal was not to identify the optimal loss function for each individual constraint (e.g., spatial relationships) or to meticulously tune their relative weights within the overall objective. Instead, we focused on demonstrating the benefits of combining these diverse constraints into a single unified cost function to improve the overall image generation process. As a result, our method may struggle with handling sophisticated instances of compositionality within each component, such as complex spatial relationships (e.g., nested spatial dependencies). This limitation, however, can be addressed in future work by replacing each component with more advanced alternatives.
Scalability Limitation. Our method is designed to perform at most two correction iterations. However, when a prompt contains a large number of entities (e.g., four or more), two iterations may not be sufficient to fully address all errors. In such cases, more than two iterations may be necessary to fully correct the output.
Appendix A Appendix
Appendix B Verifier Adaptation
@@
We modify the DA-Score (Singh & Zheng, 2023) to provide fine-grained feedback regarding the types of problems present in the image. To do so, we categorize questions by the type of issue associated with each entity. At first, similar to the DA-Score, we extract a diverse set of questions covering all entities. Then, we remove extra details from questions with one entity to create coarse-grained questions to address entity missing. Simultaneously, we categorize each question into one of the attribute binding, or spatial relation problems using LLMs (Achiam et al., 2023). After categorization, we assign a score to each question using a VQA model like DA-Score. For each entity, we aggregate question scores in every category to obtain the final fine-grained feedback. We then use the feedback provided by our evaluator for these questions to assign three scores to each entity, assessing their generation quality in each category.
Appendix C Experiments
C.1 Benchmarks
@@
HRS-Bench Benchmark. This benchmark (Bakr et al., 2023) evaluates text-to-image models across 13 skills, grouped into five key areas: accuracy, robustness, generalization, fairness, and bias. Spanning 50 scenarios, including fashion, animals, transportation, food, and clothing, it offers a comprehensive evaluation framework. To demonstrate our model’s ability to generate multi-entity images, we utilized the 3-entity prompts from its color category.
C.2 Metrics
C.2.1 Alignment Metrics
We employed various metrics to evaluate our method and compare it with other methods. The most commonly used metric for assessing Text-to-Image models is the CLIP-Score (Hessel et al., 2021), which measures the similarity between the image and text embeddings generated by the CLIP model. The TIFA-Score (Hu et al., 2023) utilizes a language model to generate multiple question-answer pairs from the input prompt, which are then filtered using a QA model. A VQA model answers these visual questions based on the generated image, and the correctness of the answers is evaluated. Additionally, the DA-Score (Singh & Zheng, 2023) breaks the prompt into a set of disjoint assertions and evaluates their alignment with the generated image directly using a VQA model. VQA models often struggle to accurately interpret spatial relationships. We employ position-based metrics as a more dependable approach for evaluating generative models with respect to spatial relationships. We utilize the VISOR Score (Gokhale et al., 2022), designed to assess the spatial relationships of entities. This score is derived by extracting entities using an object detector and calculating the positions of object bounding box centroids relative to one another. The other automated metric we used was the reward model introduced in (Xu et al., 2024). This reward model is trained on human preference comparisons to predict the alignment of generated images with textual prompts, providing an evaluation metric for the quality of text-to-image generation based on human preferences.
C.2.2 Human Evaluation
@@
C.3 More ablation Study
@@
C.3.1 Fine-grained Refinement Method Ablation Study
@@
C.3.2 Diversity & Reality
To evaluate the realism and diversity of generated images and compare models, we used the Density and Coverage metrics proposed by (Naeem et al., 2020). Density measures fidelity by assessing how many real data points lie within the neighborhood of each generated sample, reflecting how well the generated data aligns with the real distribution. Coverage quantifies diversity by determining the proportion of real data points that have at least one generated sample in their neighborhood, indicating how effectively the model captures the variability of real data.
C.4 Additional Quantitative Results
C.4.1 Attribute Binding Loss Comparison
We evaluated three different losses for attribute binding. As shown in Table 9, this analysis demonstrates the effectiveness of our IoU-based loss for addressing the attribute binding problem, outperforming other losses employed in previous studies (Li et al., 2024b).
@@ C.4.2 Additional Metrics
Similarly, as demonstrated in Table 8, our method outperforms all other models on the CLIP-Score metric for the HRS benchmark and most categories of the T2I-CompBench benchmark, while achieving comparable performance in the "Complex" and "Texture" categories. On the other hand, this metric struggles to underscore the differences between various methods, as the score variations among different models are smaller compared to other metrics. This limitation arises from the inherent weaknesses of the CLIP (Radford et al., 2021) model in compositional understanding, primarily due to its training methodology and the scoring mechanism, which overlooks several aspects of compositionally.
@@
C.5 Additional Qualitative Results
Figure 10 presents a side-by-side comparison of our method with state-of-the-art models, focusing exclusively on spatial relations, while Figure 13 compares our approach with other baselines across all categories. These figures highlight common failure cases in existing models that our method effectively addresses, such as entity missing in "a blue giraffe and a brown vase" and attribute binding errors like "a yellow banana and a red monkey." Additionally, our method successfully handles complex spatial relationships where other models struggle, as demonstrated in "a brown teddy bear holding a blue cup sits on the sofa."
Comparison of our models with fine-tuned models We compare our models with GORS (Huang et al., 2023a), a method that fine-tunes Stable Diffusion v2 using generated images closely aligned with compositional prompts. In this approach, the fine-tuning loss is weighted by a reward based on alignment scores. As demonstrated in Fig. 11, our models outperform GORS on the T2I-CompBench (Huang et al., 2023a).
@@ C.5 Additional Qualitative Results
Figure 12: Visualization of the final cross-attention maps. Compare the final cross-attention maps of Stable Diffusion v1.4 and our fine-grained initial noise refinement. The prompt is: “A black cat is on the left of a green frog.”
@@
C.6 Attention maps as object detection
We extracted object boxes from attention maps to qualitatively assess the generated images. The details for bounding box extraction are provided in section D. Fig. 8 illustrated some more visualizations of entity boxes from our generated images. Extracted boxes show great alignment with the location where the image is actually generated, showcasing the proper appearance of multiple entities.
Appendix D Implementation Details
@@
Given a caption: Specify the spatial location among objects in the caption as (object, spatial relation, object). If there is not any spatial relation in the caption, just say "No spatial" Example: Caption: "the striped rug was on top of the tiled floor" 1. (rug, on top of, floor) Example: Caption: "a couple is enjoying a picnic in the park" 1. (couple, in, park) Example: Caption: "a blue scooter is parked near a curb in front of a green vintage car" 1. (scooter, near, curb) 2. (scooter, in front of, car) Example: Caption: "the airplane is flying above the clouds." 1. (airplane, above, clouds) Example: Caption: "a bird on the left of a clock" 1. (bird, on the left of, clock) Example: Caption: "the black phone was resting on the silver charger" 1. (phone, on, charger) Example: Caption: "the bear is near a lake" 1. (bear, near, lake) Example: Caption: "a mouse on side of a bag" 1. (mouse, on side of, bag) Example: Caption: "a small white kitchen with brown wood floor" No spatial Answer as concisely as possible.
@@
Given a caption: Specify the adjective and its object in the caption as (adjective, object) If there is not any adjective in the caption, just say "No adjective" Example: Caption: "a brown dog fetching a frisbee in a green park." 1. (brown, dog) 2. (green, park) Example: Caption: "a long necklace and a short earring" 1. (long, necklace) 2. (short, earring) Example: Caption: "a soft pillow is on top of the rocking chair" 1. (soft, pillow) 2. (rocking, chair) Example: Caption: "a metallic bracelet and a leather hat" 1. (metallic, bracelet) 2. (leather, hat) Example: Caption: "a person is wearing a hat and sunglasses while fishing" No adjective Example: Caption: "a desk on the right of a horse" No adjective Example: Caption: "the shiny silver watch lay next to the smooth black leather wallet" 1. (shiny, watch) 2. (silver, watch) 3. (smooth, wallet) 4. (black, wallet) 5. (leather, wallet) Answer as concisely as possible.